OpenStudy (anonymous):

What is the equation of the line that passes through (2, 1) and is perpendicular to the graph of y = –x + 2?

OpenStudy (anonymous):

Do you remember how the slopes of two perpendicular lines are related?

OpenStudy (anonymous):

yes their negative reciprocal

OpenStudy (anonymous):

Ok, so you have all the information you need: the slope and a point(2,1)

OpenStudy (anonymous):

Y - 1 = 1 ( X -2 ) , X - Y = 1 { answer ]

OpenStudy (anonymous):

Is it clear?

OpenStudy (anonymous):

is the slope postive 1

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes, slope = -1/(-1)=1

OpenStudy (anonymous):

okay thank you thats what i was doing wrong

OpenStudy (anonymous):

You're welcome =)